I tried to solve a problem by the Elimination Method, but came up with the wrong answers. Please show me why. The following is the problem:
6x-3y=75 and
5x+7y=23
I came up with (22,19) however, the correct answers are (8,-9).
Thanks.

Answer by mananth(16949)

6 x -3 y = 75 .............1
5 x + 7 y = 23 .............2
Eliminate y
multiply (1)by 7
Multiply (2) by 3
42 x -21 y = 525
15 x 21 y = 69
Add the two equations
57 x = 594
/ 57
x = 10.42 OR 198/19
plug value of x in (1)
6 x + -3 y = 75
62.53 + -3 y = 75
-3 y = 75 -62.53
-3 y = 12.47
y = -4.16 OR -25/6
m.ananth@hotmail.ca